OpenStudy (anonymous):

Help Please? Write the equation of the circle containing the points J(-3, -3), K(1, -3), and L(1, 1)

OpenStudy (anonymous):

I'll explain :)

OpenStudy (whpalmer4):

You have 3 points, and 3 unknowns: \(h,k\) and \(r\) where \((h,k)\) is the center of the circle and \(r\) is the radius: \[(x-h)^2+(y-k)^2 = r^2\] Plug each of the three points into the formula for \(x\) and \(y\). You'll end up with three equations in three unknowns, which is what you need to solve for the three unknowns. Solve the system of equations and write the equation of the circle from the values you obtain. I'll do the first step for you: \[(x-h)^2+(y-k)^2 = r^2\]\[(-3-h)^2+(-3-k)^2 = r^2\]\[(h^2+6h+9) + (k^2+6k+9) = r^2\]\[h^2+6h+k^2+6k+18 = r^2\] Now do the other two equations in the same fashion.